OpenStudy (ggdvaaa):

The amount $5.40 is 9% of what price? $14.40 $48.60 $60.00 $97.20

OpenStudy (thephysicsman):

Set up a ratio

OpenStudy (thephysicsman):

5.49/x = 9/100 then cross multiply

jhonyy9 (jhonyy9):

5,40 ----- 9% x -------100% ------------------ x = 5,40*100/9 = ?

rebeccaxhawaii (rebeccaxhawaii):

^^^

OpenStudy (jdoe0001):

so... 5.40 is 9% of ... something anyhow, as everyone above just said :) \(\begin{array}{ccllll} value&\% \\\hline\\ 5.40&9\\ x&100 \end{array}\implies \cfrac{5.40}{x}=\cfrac{9}{100}\)

OpenStudy (ggdvaaa):

so it would beeee

OpenStudy (jdoe0001):

you tell us

OpenStudy (ggdvaaa):

A .

OpenStudy (jdoe0001):

what did "you" get?

OpenStudy (ggdvaaa):

well.

OpenStudy (ggdvaaa):

i assumed.

OpenStudy (ggdvaaa):

.

OpenStudy (jdoe0001):

just cross-multiply, as others said, and solve for "x" :)

OpenStudy (ggdvaaa):

ok

OpenStudy (ggdvaaa):

you cross mutlipy?

OpenStudy (ggdvaaa):

what exactly the example kinda confused me

OpenStudy (ggdvaaa):

ooo

OpenStudy (ggdvaaa):

its b

OpenStudy (jdoe0001):

b? how so?

OpenStudy (ggdvaaa):

it equals 0.486

OpenStudy (jdoe0001):

well, check your cross-multiplication again